OpenStudy (aaronandyson):

three consecutive natural number are such that the square of the middle number exceeds the the difference if the square of the other two by60. Assume the middle number to be x and form a quadratic equation satisfying the above statement.Hence,find the three number @sepeario

OpenStudy (sepeario):

Say the three consecutive numbers are x-1, x and x+1. Hence the equation becomes \[x^2=[(x+1)^2-(x-1)^2]+60\]

OpenStudy (sepeario):

are you able to expand the equation above?

OpenStudy (aaronandyson):

No

OpenStudy (sepeario):

ok i think you should learn how to expand these equations in your free time but for now the equation is \[x^2=(x^2+2x+1)-(x^2-2x+1)+60\] now by cancelling out what do we get?

OpenStudy (sepeario):

simplify the expression.

OpenStudy (aaronandyson):

I thought that only But unable to move any further :(

OpenStudy (aaronandyson):

Is it like x^2 - 4x - 60 = 0? x = 10 & 6?

OpenStudy (sepeario):

Very good!

OpenStudy (sepeario):

Do be aware that it is 10 and -6, which you can see from the final equation.

OpenStudy (aaronandyson):

MORE questions can u help?